OpenStudy (anonymous):

(m - 18 / 5) - (4m - 14 / 7) = 1 I need help solving this

OpenStudy (♪chibiterasu):

First we need to do the division because PEDMAS is forcing us too. >:3 18/5 = 3.6 14/7 = 2 Now we have: (m - 3.6) - (4m - 2) = 1

OpenStudy (anonymous):

Is what I have so far right? 35(m - 18 / 5) - 35(4m - 14 / 7) = 35(1) (35m - 18 / 5) - (140m - 14 / 7) = 35

OpenStudy (♪chibiterasu):

Hmm not sure where you got the 35 from. o:

OpenStudy (anonymous):

It's the LCD?

OpenStudy (♪chibiterasu):

I don't remember solving it that way. :/

OpenStudy (♪chibiterasu):

Well I guess I'll keep going: m - 3.6 - 4m - 2 = 1 We can remove the parenthesis now since none of them are like terms. Now combine like terms without the parenthesis: -3m - 5.6 = 1

OpenStudy (♪chibiterasu):

-3m - 5.6 = 1 Now add 5.6 on both sides: -3m = 6.6 And divide -3 from both sides: m = -2.2

OpenStudy (anonymous):

I don't know that way. It doesn't sound right to me..

OpenStudy (♪chibiterasu):

I think you're right, I prolly did it wrong. I keep getting 5 = 1

OpenStudy (anonymous):

Lol